# Baichuan-Omni Technical Report

### v1

@@ Abstract
The salient multimodal capabilities and interactive experience of GPT-4o highlight its critical role in practical applications, yet it lacks a high-performing open-source counterpart. In this paper, we introduce Baichuan-Omni , the first open-source 7B Multimodal Large Language Model (MLLM) adept at concurrently processing and analyzing modalities of image, video, audio, and text, while delivering an advanced multimodal interactive experience and strong performance. We propose an effective multimodal training schema starting with 7B model and proceeding through two stages of multimodal alignment and multitask fine-tuning across audio, image, video, and text modal. This approach equips the language model with the ability to handle visual and audio data effectively. Demonstrating strong performance across various omni-modal and multimodal benchmarks, we aim for this contribution to serve as a competitive baseline for the open-source community in advancing multimodal understanding and real-time interaction.
Figure 1: Evaluation across image, video, and audio modalities. (Left) Baichuan-Omni covers more modalities than Qwen2 VL [79] and outperforms the current leading omni-modal model, VITA [24]. (Right) Average scores across benchmarks for all modalities. All the scores are normalized by $x_{\text{norm}}=(x-x_{\text{min}}+10)/(x_{\text{max}}-x_{\text{min}}+10)$.
@@ 1 Introduction
In response to these challenges, we introduce an omni-modal LLM Baichuan-Omni  alongside a multimodal training scheme designed to facilitate advanced multimodal processing and naturalistic user interactions. The architecture of Baichuan-Omni is depicted in Figure 2. The scheme of Baichuan-Omni is built upon three core components:
Figure 2: Architecture of Baichuan-Omni . Baichuan-Omni is designed to process both pure text/audio inputs and combinations of video/image with text/audio. In terms of interactivity, the model initially predicts the start and end of audio inputs. During this period, incoming images and videos are encoded into features and fed into the MLLM in a streaming fashion to calculate attention. The audio features are then input into the MLLM for inference following the end of the audio input, facilitating streaming input of audio and video.
Omni-Modal Data Construction.  We utilize a substantial collection of high-quality, omni-modal data to train Baichuan-Omni with a blend of open-source, synthetic, and internally annotated datasets. In the multimodal alignment pre-training phase, we curate a wide-ranging assortment of training corpora that encompasses image captions, interleaved data, OCR data, and image-text data. For audio alignment, we collect both open-source and in-house datasets for Automatic Speech Recognition (ASR) and Audio Question Answering (AQA). In the realm of video alignment, we acquire video data from both open-source and in-house sources. During the multimodal supervised fine-tuning phase, we compile and synthesize an extensive dataset that covers over 200 tasks and comprises 600,000 instances across pure text, audio, image-text, video-text, and image-audio interaction data.
@@
Multitask Fine-tuning.  For the omni-modal fine-tuning stage, we utilize a multi-task cross-modal interaction training corpus derived from a combination of open-source, synthetic, and internally annotated data. We select data for the final supervised fine-tuning (SFT) phase based on criteria that whether factual knowledge is already learned by the pre-trained model [27]. During this phase, we implement a packing technique to concatenate multiple samples, using the cuseq_len from flash-attention2 for effective sample isolation. Through this technique, multiple samples can be packaged into a large batch while ensuring that each sample is correctly isolated during the computational process, preventing data confusion between different samples. This approach accelerates the training process and optimizes memory usage.
@@
  • We introduce Baichuan-Omni , an open-source, high-performance foundational omni-modal model capable of concurrently processing text, images, videos, and audio inputs. It also provides multilingual support for languages including English and Chinese. Our training framework features a comprehensive pipeline that includes the construction of omni-modal training data, multimodal alignment pre-training, and multimodal supervised fine-tuning, with a particular emphasis on enhancing omni-modal instruction-following capabilities.
@@
  • We have made our Baichuan-Omni model, training code, and evaluation scripts publicly available, aiming at fostering progress within the research community. As pioneers in this field, we remain committed to furthering the development of multimodal foundational models and their interactions.
@@ 2 Related works
However, compared to proprietary models like GPT-4o [63], open-source models still exhibit substantial gaps in their capabilities for multimodal interactions, and there is a considerable scarcity of open-source models that effectively facilitate comprehensive multimodal interactions [24]. To address these challenges, we propose Baichuan-Omni , an open-source and capable MLLM which concurrently supports interactions across modalities including audio, image, video, and text.
@@ 3.1 High-Quality Multimodal Data
Figure 3: Data illustration of Baichuan-Omni . We build an extensive cross-modal dataset, including text, image-text, video-text, audio-text, and their interactions. Our collection also features integrated image-audio-text and video-audio-text data.
@@ 3.2 Multimodal Alignment Pre-training
Figure 4: Training Pipeline of Baichuan-Omni . During the pretraining phase, we initially train a vision-language model using extensive image-text data, followed by training an audio-language model with ASR data. Subsequently, we integrate high-quality images, audio, and video data for comprehensive multimodal alignment. In the fine-tuning phase, we synthesize a subset of cross-modal interaction data to blend with existing high-quality datasets. From this enriched dataset, we select a subset of data that the model is already capable of handling and proceed with multimodal multitask fine-tuning. This process aims to enhance the model’s adherence to omni-modal instructions.
@@ 3.2.1 Image-Language Branch
We utilize Siglip-384px [97] as the visual encoder, which processes a 384$\times$384 image input and generates 182 tokens through a visual projector composed of a two-layer MLP and a 2$\times$2 convolution layer serving as the pooling layer. To scale the input to arbitrary resolutions while preserving the intricate details of high-resolution images, we adopt AnyRes [50], which splits the image into grids and concatenates the features of a down-sampled image to provide global context.
The training of our image-language branch is divided into three stages.
@@ 4.1.2 Major Performance
We compare Baichuan-Omni with state-of-the-art proprietary multimodal models such as Gemini 1.5 Pro [69], and GPT-4o [63], as well as a series of competitive open-source LLMs and MLLMs such as VITA [24], MAP-Neo [101], Qwen1.5-Chat [5], Llama3-Instruct [3] and OLMo [29]. We list major results on comprehensive benchmarks in Table 1.
@@
Baichuan-Omni (7B) 65.3 72.2 47.7 68.9
As shown in Table 1, our Baichuan-Omni significantly outperforms open-source, general pure-text LLMs in comprehensive benchmarks. Compared to the open-source multimodal model VITA, Baichuan-Omni demonstrates a substantial advantage in Chinese benchmarks, such as CMMLU (72.2% v.s 46.6%) and C-Eval (68.9% v.s 56.7%), and slightly surpasses VITA in AGIEval (47.7% v.s 46.2%).
@@ 4.2.1 Evaluation Benchmarks
We evaluate Baichuan-Omni on 13 representative vision-language benchmarks, including MMBench-EN, MMBench-CN [52], M3GIA [74], SEEDBench [41], RealWorldQA [87], MMMU [96], MathVista [56], MME [22], MMVet [94], TextVQA [72], OCRBench [53], ChartQA [60], and HallusionBench [30]. To ensure reproducible evaluation results, we use VLMEvalKit [20] uniformly for all evaluations. All evaluations are conducted in a zero-shot manner, adhering to the original setup of the models to ensure fair and consistent comparisons across all models and benchmarks.
@@ 4.2.2 Major Performance
We compare Baichuan-Omni with state-of-the-art proprietary multimodal models such as Gemini 1.5 Pro [69], and GPT-4o [63], as well as a series of competitive open-source multimodal models such as VITA [24] and Qwen2-VL [79]. We list major results on VQA (Visual Question Answering) benchmarks and results on MCQ (Multi-choice & Yes-or-No Question) benchmarks in Table 2 and Table 3.
@@
Baichuan-Omni (7B) 62.6 65.4 51.9 74.3 79.6 70.0
As shown in Table 2 and Table 3, our Baichuan-Omni comprehensively outperformed VITA-8*7b [24], which has 12B activated parameters, in multiple visual tasks, both on VQA benchmarks and MCQ benchmarks. Besides, we also demonstrates competitive performance comparable to, or even better than, open-source image-specialized models like MiniCPM-Llama3-V 2.5 [92]. Specifically, Baichuan-Omni outperformed MiniCPM-Llama3-V 2.5 on most VQA tasks, including MMBench-CN, SEED-IMG, MME, HallusionBnech and MMMU which requres expert-level perception and reasoning. However, despite the advantage of incorporating an additional audio modality compared to Qwen2-VL [79], the performance gap between our model and Qwen2-VL in image tasks remains evident. Furthermore, it is worth noting that, beyond Qwen2-VL, the stark divide between open-source and closed-source models remains substantial.
@@ 4.3.1 Evaluation Benchmarks
We perform a thorough evaluation on general video understanding tasks (General VQA) and open-ended video question answering (Open-ended VQA) to comprehensively assess the video understanding capabilities of Baichuan-Omni .
@@ 4.3.2 Major Performance
We compare Baichuan-Omni with state-of-the-art multimodal proprietary models such as Gemini 1.5 Pro [69], GPT 4V [62], and GPT 4o [63], and a series of competitive open-source multimodal models such as VITA [24], Qwen2-VL [79], AnyGPT [98], VideoLLaMA 2 [13], VideoChat2 [44], LLaVA-NeXT-Video [104], and Video-LLaVA [48]. We list major results on general video understanding benchmarks in Table 4 and results on open-ended video question answering in Table 5.
@@
Baichuan-Omni (7B) 1 fps (max 48) 60.9 58.8 58.2 56.8
Results on general video understanding benchmarks. As shown in Table 4, Baichuan-Omni demonstrates competitive results over proprietary models on benchmarks like Egoschema and MVBench, and achieves strong performance across open-source multimodal models, which shows comprehensive video understanding capabilities of Baichuan-Omni .
Compared to VITA, a MoE omni-modal LLM with about 12B activated parameters, Baichuan-Omni (7B) outperforms it on all General Video QA benchmarks, and achieve an average improvement of about 4%. Additionally, Baichuan-Omni excels a series of open-source models such as VideoLLaMA 2, VideoChat2, LLaVA-NeXT-Vide, and Video-LLaVA. Notably, Baichuan-Omni also outperforms the proprietary model GPT 4V on MVBench (43.7%) and Egoschema (55.6%).
Results on open-ended video question answering benchmarks. The performance on Open-ended VQA is listed in Table 5. Baichuan-Omni demonstrates SoTA performance (both Accuracy and Score) on ActivityNet-QA and MSVD-QA across all open-source models, such as the most recent SoTA multimodal models VITA and Qwen2 VL, and outperforms the proprietary model Gemini 1.5 Pro (56.7%) on ActivityNet-QA. The superior results indicate that Baichuan-Omni is also effective in open-ended question answering, i.e., Baichuan-Omni is more capable of generating informative and descriptive responses.
@@ 4.4.1 Evaluation Benchmarks
To validate the audio understanding capacity of Baichuan-Omni , we present the evaluating results on benchmarks with three tasks:
@@ 4.4.2 Major Performance
Baichuan-Omni is compared with the state-of-the-art baselines across ASR, S2TT and SER tasks, including the recent leading large audio-language model Qwen2-Audio-Instruct [14] and the large omni-modal language model VITA [24]. On top of that, the performance of the classical pre-trained audio language model, Whisper-large-v3 [67], is presented for ASR and the performance of SALMONN [77] is presented for S2TT.
Results on ASR benchmarks. Baichuan-Omni exhibits a strong audio transcription capacity in Table 6. Baichuan-Omni primarily targets the Chinese corpus. In the general Chinese ASR scene, Baichuan-Omni has a 2.0% WER (2.6% CER) superiority on the Fleurs test-zh subset and 4.1% WER (4.2% CER) improvement on the WenetSpeech test_net when comparing with Qwen2-Audio-Instruct. The evaluation results on WenetSpeech consistently demonstrate the superiority of our model over VITA. Baichuan-Omni achieves nearly a 50% improvement in the CER performance of VITA, both in test_net (7.1% v.s 12.2%) and test_meeting (8.9% v.s 16.5%) subsets. On the more challenging Chinese dialect benchmark, KeSpeech, our model maintains a comprehensive lead, with an average CER of 6.7% over the performance of all dialects. Notably, while our model excels in Chinese audio transcription, Baichuan-Omni also maintains robust general ASR performance in English. We achieve 4.7% of WER, which exceeds Qwen2-Audio-Instruct by 11% WER.
@@
Results on S2TT and AIR-Bench benchmarks. In addition to ASR, Baichuan-Omni excels in both S2TT and SER tasks. The evaluation results are summarized in Table 7. Notably, when translating from English to Chinese on the Covost-2 en2zh test set, Baichuan-Omni outperforms Qwen2-Audio-Instruct by approximately 7 BLEU points. For the reverse translation, from Chinese to English, our model’s performance on the Covost-2 zh2en test set is comparable to that of Qwen2-Audio-Instruct. On the AirBench, Baichuan-Omni scores 7.42 and 7.26 for speech and sound, respectively, outperforming Qwen2-Audio-Instruct and showcasing Baichuan-Omni ’s superior ability to generate realistic human speeches and sounds.

### v3

# Ocean-omni: To Understand the World with Omni-modality

@@ Abstract
The salient multimodal capabilities and interactive experience of GPT-4o highlight its critical role in practical applications, yet it lacks a high-performing open-source counterpart. In this paper, we introduce Ocean-omni , the first open-source 7B Multimodal Large Language Model (MLLM) adept at concurrently processing and analyzing modalities of image, video, audio, and text, while delivering an advanced multimodal interactive experience and strong performance. We propose an effective multimodal training schema starting with 7B model and proceeding through two stages of multimodal alignment and multitask fine-tuning across audio, image, video, and text modal. This approach equips the language model with the ability to handle visual and audio data effectively. Demonstrating strong performance across various omni-modal and multimodal benchmarks, we aim for this contribution to serve as a competitive baseline for the open-source community in advancing multimodal understanding and real-time interaction.
Figure 1: Evaluation across image, video, and audio modalities. (Left) Ocean-omni covers more modalities than Qwen2 VL [79] and outperforms the current leading omni-modal model, VITA [24]. (Right) Average scores across benchmarks for all modalities. All the scores are normalized by $x_{\text{norm}}=(x-x_{\text{min}}+10)/(x_{\text{max}}-x_{\text{min}}+10)$.
@@ 1 Introduction
In response to these challenges, we introduce an omni-modal LLM Ocean-omni  alongside a multimodal training scheme designed to facilitate advanced multimodal processing and naturalistic user interactions. The architecture of Ocean-omni is depicted in Figure 2. The scheme of Ocean-omni is built upon three core components:
Figure 2: Architecture of Ocean-omni . Ocean-omni is designed to process both pure text/audio inputs and combinations of video/image with text/audio. In terms of interactivity, the model initially predicts the start and end of audio inputs. During this period, incoming images and videos are encoded into features and fed into the MLLM in a streaming fashion to calculate attention. The audio features are then input into the MLLM for inference following the end of the audio input, facilitating streaming input of audio and video.
Omni-Modal Data Construction.  We utilize a substantial collection of high-quality, omni-modal data to train Ocean-omni with a blend of open-source, synthetic, and internally annotated datasets. In the multimodal alignment pre-training phase, we curate a wide-ranging assortment of training corpora that encompasses image captions, interleaved data, OCR data, and image-text data. For audio alignment, we collect both open-source and in-house datasets for Automatic Speech Recognition (ASR) and Audio Question Answering (AQA). In the realm of video alignment, we acquire video data from both open-source and in-house sources. During the multimodal supervised fine-tuning phase, we compile and synthesize an extensive dataset that covers over 200 tasks and comprises 600,000 instances across pure text, audio, image-text, video-text, and image-audio interaction data.
@@
Multitask Fine-tuning.  For the omni-modal fine-tuning stage, we utilize a multi-task cross-modal interaction training corpus derived from a combination of open-source, synthetic, and internally annotated data. We select data for the final supervised fine-tuning (SFT) phase based on criteria that whether factual knowledge is already learned by the pre-trained model [27]. During this phase, we implement a packing technique to concatenate multiple samples, using the cuseq_len from flash-attention2 for effective sample isolation. With this, multiple samples can be packaged into a large batch while ensuring each sample is correctly isolated during the computational process, preventing data confusion between different samples. This accelerates the training process and optimizes memory usage.
@@
  • We introduce Ocean-omni , an open-source, high-performance foundational omni-modal model capable of concurrently processing text, images, videos, and audio inputs. It also provides multilingual support for languages including English and Chinese. Our training framework features a comprehensive pipeline that includes the construction of omni-modal training data, multimodal alignment pre-training, and multimodal supervised fine-tuning, with a particular emphasis on enhancing omni-modal instruction-following capabilities.
@@
  • We have made our Ocean-omni model, training code, and evaluation scripts publicly available, aiming at fostering progress within the research community. As pioneers in this field, we remain committed to furthering the development of multimodal foundational models and their interactions.
@@ 2 Related works
However, compared to proprietary models like GPT-4o [63], open-source models still exhibit substantial gaps in their capabilities for multimodal interactions, and there is a considerable scarcity of open-source models that effectively facilitate comprehensive multimodal interactions [24]. To address these, we propose Ocean-omni , an open-source and capable MLLM which concurrently supports interactions across modalities including audio, image, video, and text.
@@ 3.1 High-Quality Multimodal Data
Figure 3: Data illustration of Ocean-omni . We build an extensive cross-modal dataset, including text, image-text, video-text, audio-text, and their interactions. Our collection also features integrated image-audio-text and video-audio-text data.
@@ 3.2 Multimodal Alignment Pre-training
Figure 4: Training Pipeline of Ocean-omni . During the pretraining phase, we initially train a vision-language model using extensive image-text data, followed by training an audio-language model with ASR data. Subsequently, we integrate high-quality images, audio, and video data for comprehensive multimodal alignment. In the fine-tuning phase, we synthesize a subset of cross-modal interaction data to blend with existing high-quality datasets. From this enriched dataset, we select a subset of data that the model is already capable of handling and proceed with multimodal multitask fine-tuning. This process aims to enhance the model’s adherence to omni-modal instructions.
@@ 3.2.1 Image-Language Branch
We utilize Siglip-384px [97] as the visual encoder, which processes a 384$\times$384 image input and generates 182 tokens through a visual projector composed of a two-layer MLP and a 2$\times$2 convolution layer serving as the pooling layer. To scale the input to arbitrary resolutions while preserving the intricate details of high-resolution images, we adopt AnyRes [50], which splits the image into grids and concatenates the features of a down-sampled image to provide global context. The training of our image-language branch is divided into three stages.
@@ 4.1.2 Major Performance
We compare Ocean-omni with state-of-the-art proprietary multimodal models such as Gemini 1.5 Pro [69], and GPT-4o [63], as well as a series of competitive open-source LLMs and MLLMs such as VITA [24], MAP-Neo [101], Qwen1.5-Chat [5], Llama3-Instruct [3] and OLMo [29]. We list major results on comprehensive benchmarks in Table 1.
@@
Ocean-omni (7B) 65.3 72.2 47.7 68.9
As shown in Table 1, our Ocean-omni significantly outperforms open-source, general pure-text LLMs in comprehensive benchmarks. Compared to the open-source multimodal model VITA, Ocean-omni demonstrates a substantial advantage in Chinese benchmarks, such as CMMLU (72.2% v.s 46.6%) and C-Eval (68.9% v.s 56.7%), and slightly surpasses VITA in AGIEval (47.7% v.s 46.2%).
@@ 4.2.1 Evaluation Benchmarks
We evaluate Ocean-omni on 13 representative vision-language benchmarks, including MMBench-EN, MMBench-CN [52], M3GIA [74], SEEDBench [41], RealWorldQA [87], MMMU [96], MathVista [56], MME [22], MMVet [94], TextVQA [72], OCRBench [53], ChartQA [60], and HallusionBench [30]. To ensure reproducible evaluation results, we use VLMEvalKit [20] uniformly for all evaluations. All evaluations are conducted in a zero-shot manner, adhering to the original setup of the models to ensure fair and consistent comparisons across all models and benchmarks.
@@ 4.2.2 Major Performance
We compare Ocean-omni with state-of-the-art proprietary multimodal models such as Gemini 1.5 Pro [69], and GPT-4o [63], as well as a series of competitive open-source multimodal models such as VITA [24] and Qwen2-VL [79]. We list major results on VQA (Visual Question Answering) benchmarks and results on MCQ (Multi-choice & Yes-or-No Question) benchmarks in Table 2 and Table 3.
@@
Ocean-omni (7B) 62.6 65.4 51.9 74.3 79.6 70.0
As shown in Table 2 and Table 3, our Ocean-omni comprehensively outperformed VITA-8*7b [24], which has 12B activated parameters, in multiple visual tasks, both on VQA benchmarks and MCQ benchmarks. Besides, we also demonstrates competitive performance comparable to, or even better than, open-source image-specialized models like MiniCPM-Llama3-V 2.5 [92]. Specifically, Ocean-omni outperformed MiniCPM-Llama3-V 2.5 on most VQA tasks, including MMBench-CN, SEED-IMG, MME, HallusionBnech and MMMU which requres expert-level perception and reasoning. However, despite the advantage of incorporating an additional audio modality compared to Qwen2-VL [79], the performance gap between our model and Qwen2-VL in image tasks remains evident. Furthermore, it is worth noting that, beyond Qwen2-VL, the stark divide between open-source and closed-source models remains substantial.
@@ 4.3.1 Evaluation Benchmarks
We perform a thorough evaluation on general video understanding tasks (General VQA) and open-ended video question answering (Open-ended VQA) to comprehensively assess the video understanding capabilities of Ocean-omni .
@@ 4.3.2 Major Performance
We compare Ocean-omni with state-of-the-art multimodal proprietary models such as Gemini 1.5 Pro [69], GPT 4V [62], and GPT 4o [63], and a series of competitive open-source multimodal models such as VITA [24], Qwen2-VL [79], AnyGPT [98], VideoLLaMA 2 [13], VideoChat2 [44], LLaVA-NeXT-Video [104], and Video-LLaVA [48]. We list major results on general video understanding benchmarks in Table 4 and results on open-ended video question answering in Table 5.
@@
Ocean-omni (7B) 1 fps (max 48) 60.9 58.8 58.2 56.8
Results on general video understanding benchmarks. As shown in Table 4, Ocean-omni demonstrates competitive results over proprietary models on benchmarks like Egoschema and MVBench, and achieves strong performance across open-source multimodal models, which shows comprehensive video understanding capabilities of Ocean-omni .
Compared to VITA, a MoE omni-modal LLM with about 12B activated parameters, Ocean-omni (7B) outperforms it on all General Video QA benchmarks, and achieve an average improvement of about 4%. Additionally, Ocean-omni excels a series of open-source models such as VideoLLaMA 2, VideoChat2, LLaVA-NeXT-Vide, and Video-LLaVA. Notably, Ocean-omni also outperforms the proprietary model GPT 4V on MVBench (43.7%) and Egoschema (55.6%).
Results on open-ended video question answering benchmarks. The performance on Open-ended VQA is listed in Table 5. Ocean-omni demonstrates SoTA performance (both Accuracy and Score) on ActivityNet-QA and MSVD-QA across all open-source models, such as the most recent SoTA multimodal models VITA and Qwen2 VL, and outperforms the proprietary model Gemini 1.5 Pro (56.7%) on ActivityNet-QA. The superior results indicate that Ocean-omni is also effective in open-ended question answering, i.e., Ocean-omni is more capable of generating informative and descriptive responses.
@@ 4.4.1 Evaluation Benchmarks
To validate the audio understanding capacity of Ocean-omni , we present the evaluating results on benchmarks with three tasks:
@@ 4.4.2 Major Performance
Ocean-omni is compared with the state-of-the-art baselines across ASR, S2TT and SER tasks, including the recent leading large audio-language model Qwen2-Audio-Instruct [14] and the large omni-modal language model VITA [24]. On top of that, the performance of the classical pre-trained audio language model, Whisper-large-v3 [67], is presented for ASR and the performance of SALMONN [77] is presented for S2TT.
Results on ASR benchmarks. Ocean-omni exhibits a strong audio transcription capacity in Table 6. Ocean-omni primarily targets the Chinese corpus. In the general Chinese ASR scene, Ocean-omni has a 2.0% WER (2.6% CER) superiority on the Fleurs test-zh subset and 4.1% WER (4.2% CER) improvement on the WenetSpeech test_net when comparing with Qwen2-Audio-Instruct. The evaluation results on WenetSpeech consistently demonstrate the superiority of our model over VITA. Ocean-omni achieves nearly a 50% improvement in the CER performance of VITA, both in test_net (7.1% v.s 12.2%) and test_meeting (8.9% v.s 16.5%) subsets. On the more challenging Chinese dialect benchmark, KeSpeech, our model maintains a comprehensive lead, with an average CER of 6.7% over the performance of all dialects. Notably, while our model excels in Chinese audio transcription, Ocean-omni also maintains robust general ASR performance in English. We achieve 4.7% of WER, which exceeds Qwen2-Audio-Instruct by 11% WER.
@@
Results on S2TT and AIR-Bench benchmarks. In addition to ASR, Ocean-omni excels in both S2TT and SER tasks. The evaluation results are summarized in Table 7. Notably, when translating from English to Chinese on the Covost-2 en2zh test set, Ocean-omni outperforms Qwen2-Audio-Instruct by approximately 7 BLEU points. For the reverse translation, from Chinese to English, our model’s performance on the Covost-2 zh2en test set is comparable to that of Qwen2-Audio-Instruct. On the AirBench, Ocean-omni scores 7.42 and 7.26 for speech and sound, respectively, outperforming Qwen2-Audio-Instruct and showcasing Ocean-omni ’s superior ability to generate realistic human speeches and sounds.